A clinical trial tests a method designed to increase the probability of conceiving a girl. In the study 480 babies were born, and 264 of them were girls. Use the sample data to construct a 99% confidence interval estimate of the percentage of girls born. Based on the result, does the method appear to be effective?
__
(Please show all work, thank you.)

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

A clinical trial tests a method designed to increase the probability of conceiving a girl. In the study 480 babies were born, and 264 of them were girls. Use the sample data to construct a 99% confidence interval estimate of the percentage of girls born. Based on the result, does the method appear to be effective?
-----
sample proportion = p-hat = 264/480 = 0.55
ME = 2.5758*sqrt[0.55*0.45/480] = 0.0585
----------------------
99% CI: 0.55-0.0585 < p < 0.55+0.0585
------
0.49< p < 0.55
===================
Conclusion:: You be the judge.
======================
Cheers,
Stan H.
----------------------

Answer by ewatrrr(24785)

Ho: u = .50 (probability of having a girl w/o treatment)
Ha: u > .50 for clinical trial
mean = .55
SD = sqrt(.55*.45) = .4975
t = (.55-.50)/(.4975/sqrt(480) = .05/.0227 = 2.202
P(t > 2.202) = .0141
At the 99% confidence interval, .0141 > .01
there is not enough evidence to indicate method is working